# Memory-less and Backscatter-less Tunnel Diode Harmonic Signatures for RFID

### v1

@@ 5 N/A
TABLE V: Predicted Fundamental Frequencies (Wireless) at 200 mV from the Wireless Measurements
@@ V-C Time Domain Measurements
Using an oscilloscope, the wireless outputs from the tunnel diode boards were captured and measured when the boards were producing relatively stable harmonics and when they were not. This work demonstrates the transient behavior of Board 1 at 139 and 201 mV in Fig. 14 and Fig. 15, respectively. Shown in Fig. 15, at 201 mV, Board 1 produced a relatively distinct sinusoid, with the fast-Fourier-transform (FFT) analysis from the oscilloscope, showing also distinct harmonics. This result is consistent with Board 1’s behavior at 200 mV in Fig. 6. At 139 mV, Board 1 was entering the negative differential resistance region, and the output of the board was ”bursty,” as shown in the left-most plot of Fig. 14. Within each burst is a Gaussian pulse-like shape, hence the spread-out spectrum FFT results in the right most plot of Fig. 14. Similarly, Board 5 at 200 mV, in Fig. 16, demonstrated a Gaussian pulse-like output and also a spread-out FFT result. Again, this behavior is consistent with what was demonstrated by Board 5 in Fig. 6. Interestingly, these bursts that Board 1 at 139 mV and Board 5 at 200 mV demonstrated suggest that the tunnel diode oscillator boards act as self-blocking or even super-regenerative oscillators [19] as the measured output ramps up and then decays to build up oscillation for a new burst.
VI Link Budget
@@
Figure 17: Theoretical link budget for tunnel diode Board 1.
@@
Using the over-the-cable measurement results and the power transmitted by each harmonic ($P_{T}$) for Board 1 at 200 mV, a theoretical link budget between the tunnel diode board and a receiving RFID reader is calculated and shown in Fig. 17. The purpose of this link budget is to estimate the range of harmonic detection and determine specific design parameters for future reader design to improve sensing/detection range.
@@
In this link budget, the antennas for both the board and reader are assumed to be 3 dBi flat over a wide bandwidth, and the reader sensitivity is set to match the minimum sensitivity of the Impinj R700 reader at -90 dBm [20]. As shown by doing this link budget, the fundamental (first) and second harmonics can be detected by a reader of that sensitivity past 50 meters. The third and fourth harmonic would not be detected by the reader for more than about 5.5 meters.
An important factor, indicative by this link budget, is to have a wide-band antenna. One of higher gain can help improve the power received at the reader. Additionally, the reader must be designed to be wide-band and have low sensitivity to detect the higher order harmonics.
@@ VII Harmonic Signatures and Applications
Tunnel diodes can be included in future tag designs as part of their backscatter modulator circuitry. To actuate the generation of harmonic signatures from these diodes, semi-passive tags can periodically bias these tunnel diodes to wake-up and emit a harmonic signature, similar to Bluetooth beacons’ advertisements. Passive tags can also bias these tunnel diodes when seeing a ”wake-up” signal from the reader, possibly without fully powering on the whole tag’s circuitry. These tags’ tunnel diodes would only need to be biased in their negative differential resistance regions to produce distinct signatures. And based on the board designs presented in this work, these signatures can be of high quality factor as those exhibited by the AI101E diodes. There are two possible useful applications for having RFID reader receiving harmonic signatures from tunnel diodes: detecting RFID tags’ presence in an area and identifying tags.
@@ VII-A Tag Detection
The simplest use of harmonics is to aid RFID readers in determining if there are tags present in an area and even attempt to localize them. Using harmonics for detection have originated before the year 2000 with the use of ”harmonic radar” and non-linear tags to track beetles [21] and bees [22], since insect-tracking requires mitigating the effects of strong clutter interference. Similarly, a RFID reader operates in an environment filled with clutter, other readers and transmitters, and strong self-interference. A reader can discover tags with their emitted harmonics and then go through with the standard communication protocols with the tags. The benefit of using harmonics for sensing tags, versus just listening for the tag’s backscatter in the industrial-scientific-medicine (ISM) band, is that it mitigates self-interference issues as the harmonics are located away from the carrier-frequency in that band and loosens the reader’s receive chain frequency cut-off sharpness when filtering out self-interference. Additionally, the use of harmonics that lie away from the carrier frequency band allows for improved localization of tags as demonstrated by Ma and Kan [23] and Hui et. al [24].
@@ VIII Conclusion
This work presented the characterization of tunnel diodes, behaving as a harmonics-producing oscillator and demonstrated these diodes’ capability to produce hardware-intrinsic ”memory-less” and ”backscatter-less” harmonic signatures for identification. Future studies, based on this current work, should investigate optimal tag antenna design to allow for wide-band sensing of these diodes’ harmonic signatures wireless. A big limiting factor for using tunnel diodes for harmonic RFID would be the tag’s antenna. Traditionally, a narrow-band antenna acts as a ”filter” that dampens out-of-band harmonics from being backscattered. As shown by the spacing of the harmonics in Fig. 6, a wide-band antenna is needed by both the tag and reader to transmit and capture the almost random placement of these harmonics. If not a wide-band antenna, then an antenna that periodically resonates at a specific tunnel diode’s harmonics can also be used. Future wide-band reader designs also should be investigated for processing of wide-band harmonic signatures. Similar to other wide-band receivers, a harmonic RFID reader will have to deal with multiple non-RFID tag interfering transmitters that can desensitize the reader. Additional future wireless experimentation of tunnel diode tags should investigate and measure sensing ranges for these tags and sensing and communicating with a group of these tags. For practicalness, methods to ensure harmonic tunnel diode RFID tags meet the regulations on ultra-wide-band wireless devices should also be investigated.
@@ IX Acknowledgments
The authors of this paper wants to thank Walter (Drew) Disharoon for aiding with the wireless measurements and Dr. Nima Ghalichechian and his mmWave Antennas and Arrays Laboratory for providing the anechoic chamber, antennas, and other supporting equipment for the wireless measurements. Thank you also to Zachary Brumbach and Dr. John Cressler for providing the spectrum analyzer for the wireless measurements.

### v2

@@ 5 N/A
TABLE VI: Predicted Fundamental Frequencies (Wireless) at 200 mV from the Wireless Measurements
@@ V-C Time Domain Measurements
Using an oscilloscope, the wireless outputs from the tunnel diode boards were captured and measured when the boards were producing relatively stable harmonics and when they were not. This work demonstrates the transient behavior of Board 1 at 139 mV and 201 mV in Fig. 16 and Fig. 17, respectively. Shown in Fig. 17, at 201 mV, Board 1 produced a relatively distinct sinusoid, with the Fast-Fourier-Transform (FFT) analysis from the oscilloscope, showing also distinct harmonics. This result is consistent with Board 1’s behavior at 200 mV in Fig. 7. At 139 mV, Board 1 was entering the negative differential resistance region, and the output of the board was ”bursty,” as shown in the left-most plot of Fig. 16. Within each burst is a Gaussian pulse-like shape, hence the spread-out spectrum FFT results in the right most plot of Fig. 16. Similarly, Board 5 at 200 mV, in Fig. 18, demonstrated a Gaussian pulse-like output and also a spread-out FFT result. Again, this behavior is consistent with what was demonstrated by Board 5 in Fig. 7. Interestingly, these bursts that Board 1 at 139 mV and Board 5 at 200 mV demonstrated suggest that the tunnel diode oscillator boards act as self-blocking or even super-regenerative oscillators [25] as the measured output ramps up and then decays to build up oscillation for a new burst.
VI Link Budgets and Reader and Harmonic Tag Hardware Design Considerations
A wide-band, harmonic RFID reader and tag with a harmonic tunnel diode oscillator have to be designed with special considerations in regards to power consumption of the tag, read-range, and frequency planning for wide-band operation. Preliminary link-budgets of the forward and reverse links, with tunnel diode Board 1 serving as the tag, are presented to help theorize possible ranges and provide insight on design choices for the reader and tag. The reverse and forward links are significant since they are related to read-range and power-up range, respectively.
VI-A Reverse Link
@@
Figure 19: Theoretical reverse link budget for tunnel diode Board 1.
@@
The reverse link range is the range the information from a harmonic tag can be read by the reader’s receiver. Using the over-the-cable measurement results and the power transmitted ($P_{T}$) by each harmonic for Board 1 at 200 mV, a theoretical one-way link budget between the tunnel diode board, acting as a tag, and a receiving RFID reader (RX) is calculated and shown in Fig. 19. The purpose of this link budget is to estimate the range of harmonic detection and determine specific design parameters for future reader receiver design to improve sensing/detection range. Note, the transmitted signal strength of the reverse link (from tag to reader) is not directly affected by the forward link (from tag to reader) since no backscattering is involved.
@@
In this link budget, the antennas for both the board and reader are assumed to be 3 dBi flat over the wide bandwidth that includes the fundamental and harmonic frequencies of Board 1. The reader sensitivity is set and proposed to match the rated sensitivity or lowest detection level of the Impinj R700 reader at -90 dBm [26]. As shown by doing this link budget, the fundamental (first) and second harmonics can be detected by a reader of that sensitivity past 50 meters. The third and fourth harmonic would not be detected by the reader for more than about 5.5 meters. Better reader sensitivity is needed to detect these later harmonics past 5 meters. Another important factor, indicative by this link budget, is to have a wide-band antenna. One of higher gain can help improve the power received at the reader. The reader must be designed to be wide-band and have low minimum-power detection sensitivity level to detect the higher order harmonics.
VI-B Forward Link
Figure 20: Theoretical forward link budget for tunnel diode board 1.
[…]
As demonstrated by these results, a fully-passive design would be forward-range limited. At 200 mV, Board 1’s tunnel diode consumed 524.6 ($\mu W$) of power, which is significantly higher than the tunnel diode power consumption shown in [3], which consumed 20 $\mu W$. Thus, using the higher power-consuming tunnel diode would only allow the tag to be powered about 2.74 meters compared to 14.27 meters when using a lower-powered tunnel diode, such at the 20 $\mu W$ tunnel diode used in [3]. The forward-range limitations can be resolved by using a semi-passive design that only biases the diode when the tag’s ID is requested.
VI-C System Hardware Design Considerations
The main issues for this theorized harmonic RFID system are: diode’s power-consumption and reader’s sensitivity. As shown by the link budgets for the forward and reverse links, the system is more constrained with the forward-range compared to reverse-range due to the power-consumption of the tunnel diode. The tag would have better forward-range if it is semi-passive. Then, the reader will transmit a ”wake-up” signal, and the tag can utilize a simple, passive wake-up receiver to then detect the wake-up signal and only bias the tunnel diode to emit a harmonic identifying signal without powering on the other components and processes of the tag.
The reader can be designed as traditional up-conversion transmit and down-conversion receive chains, with looser power and size constraints than the tag. Hence, the reader can use traditional mixers, amplifiers, and other active components. For improved receiver sensitivity, careful planning of the RF chain needs to be done to avoid desensitization by unwanted signals and only amplification of the wanted harmonic content, which can be challenging for wide-band ($>$500 MHz) operation. For the reverse-range, since the harmonic signature is wide-band, then the reader’s receiver hardware will have to operate wide-band and need careful frequency planning to mitigate or filter out blockers or interferers that can saturate and desensitize the hardware. The wide-band operation also provides challenges in sourcing or designing the traditional RF components of a receive chain (such as amplifier, filters, etc.) to have strong performance and a flat response over a wide-band. Additionally, the antenna design for both the tag and reader’s receiver needs to be designed to be wide-band to be able to transmit and capture the widely-spaced out harmonics from the tag.
@@ VII Harmonic Signatures and Applications
Tunnel diodes can be included in future tag designs as part of their backscatter modulator circuitry. To actuate the generation of harmonic signatures from these diodes, semi-passive tags can periodically bias these tunnel diodes to wake-up and emit a harmonic signature, similar to Bluetooth beacons’ advertisements. These tags’ tunnel diodes would only need to be biased in their negative differential resistance regions to produce distinct signatures. Shown by the measurements presented in this work, these signatures can be of high quality factor as those exhibited by the AI101E diodes. There are two possible useful applications for having RFID reader receiving harmonic signatures from tunnel diodes: detecting RFID tags’ presence in an area and identifying tags.
@@ VII-A Tag Detection
The simplest use of harmonics is to aid RFID readers in determining if there are tags present in an area and even attempt to localize them. Using harmonics for detection have originated before the year 2000 with the use of ”harmonic radar” and non-linear tags to track beetles [28] and bees [29], since insect-tracking requires mitigating the effects of strong clutter interference. Similarly, a RFID reader operates in an environment filled with clutter, other readers and transmitters, and strong self-interference. A reader can discover tags with their emitted harmonics and then go through with the standard communication protocols with the tags. The benefit of using harmonics for sensing tags, versus just listening for the tag’s backscatter in the industrial-scientific-medicine (ISM) band, is that it mitigates self-interference issues as the harmonics are located away from the carrier-frequency in that band and loosens the need for the reader’s receive chain to have sharp filter cut-offs when filtering out self-interference. Additionally, the use of harmonics that lie away from the carrier frequency band allows for improved localization of tags as demonstrated by Ma and Kan [30] and Hui et. al [31].
@@ VII-B Tag Identification
For a harmonic RFID reader, its receiver can perform several techniques to read the harmonic identifications from tags with tunnel diode oscillators. A software-defined receiver would be able to perform a FFT and thresholding. Or like a frequency-shift-key (FSK) detection scheme, the reader can perform correlation or matched filtering and envelope-detection [33] to detect known signatures. The process of learning and later identifying and verifying these harmonic signatures lead into the concept of RFID fingerprinting.
@@ VIII Conclusion
This work presented the characterization of tunnel diodes, behaving as a harmonics-producing oscillator and demonstrated these diodes’ capability to produce hardware-intrinsic ”memory-less” and ”backscatter-less” harmonic signatures for identification. Future studies, based on this current work, should investigate optimal tag antenna design to allow for wide-band sensing of these diodes’ harmonic signatures wireless. A big limiting factor for using tunnel diodes for harmonic RFID would be the tag’s antenna. Traditionally, a narrow-band antenna acts as a ”filter” that dampens out-of-band harmonics from being emitted. As shown by the spacing of the harmonics in Fig. 7, a wide-band antenna is needed by both the tag and reader to transmit and capture the almost random placement of these harmonics, respectively. If not a wide-band antenna, then an antenna that periodically resonates at a specific tunnel diode’s harmonics can also be used. Future wide-band reader designs, with high sensitivity for detecting weaker higher-order harmonics, also should be investigated for detecting and processing wide-band harmonic signatures. Similar to other wide-band receivers, a harmonic RFID reader will have to deal with multiple non-RFID tag interfering transmitters that can desensitize the reader and need wide-band components, which bring challenges to the reader’s design and frequency planning. Additional future wireless experimentation of tunnel diode tags should investigate and measure realistic sensing ranges for these tags and possible protocols for sensing and communicating with a group of these tags. For practicalness, methods to ensure harmonic tunnel diode RFID tags meet the regulations on ultra-wide-band wireless devices should also be investigated.
@@ IX Acknowledgments
The authors of this paper want to thank Walter (Drew) Disharoon for aiding with the wireless measurements and Dr. Nima Ghalichechian and his mmWave Antennas and Arrays Laboratory for providing the anechoic chamber, antennas, and other supporting equipment for the wireless measurements. Thank you also to Zachary Brumbach and Dr. John Cressler for providing the spectrum analyzer for the wireless measurements.